Suppose a bicycle tire is inflated to a pressure of 70.0 kPa at 25 °C. If the bicycle is taken outside, where the temperature is -3 °C, what is the new pressure of the gas inside the tire? Assume the volume and amount of gas remain constant. Show your work and explain each step


1
Expert's answer
2021-01-21T03:50:12-0500

"P_1 = 70.0 \;kPa \\

T_1 = 25 + 273 = 298 \;K \\

T_2 = -3 + 273 = 270 \;K \\

V = constant \\

\frac{P_1}{T_1} = \frac{P_2}{T_2} \\

P_2 = \frac{P_1 \times T_2}{T_1} \\

= \frac{70.0 \times 270}{298} \\

= 68.42 \;kPa"

Answer: 68.42 kPa